How many grams of NaF would have to be added to 2.00 L of 0.100 M HF

to yield a solution with a pH = 4.00?


1
Expert's answer
2022-04-26T14:28:02-0400

57.12 g if Ka (HF) = 6.8×10-4

56.28 g if Ka (HF) = 6.7×10-4

60.48 g if Ka (HF) = 7.2×10-4